# Global Bifurcation of Non-Radial Solutions for Symmetric Sub-linear Elliptic Systems on the Planar Unit Disc

### v2

@@ 1. Introduction
We draw our motivation from natural phenomena that are modeled by parameterized systems of autonomous Partial Differential Equations (PDEs), where the possible states of each phenomenon are represented by solutions of the associated system of equations at corresponding parameter values. These solutions can be categorized as trivial or non-trivial. Trivial solutions are so-named because they persist for all parameter values and their existence is straightforward to determine. In contrast, non-trivial solutions may only exist for a particular range of parameter values and often exhibit exotic properties that enhance our understanding of the model. The classical bifurcation problem is concerned with the existence and global behaviour of branches of non-trivial solutions emerging from their trivial counterparts. Phenomena which admit the symmetries of a certain group $G$ (represented by the $G$-equivariance of the associated equations) may be studied as equivariant bifurcation problems, with additional consideration placed on the symmetric properties of these branches (cf. [1, 3, 9, 11, 19, 21, 22, 23, 24, 25]). The application of topological methods to the study of differential equations, as with most tools in the arsenal of the nonlinear analyst, traces back to Maître Poincare [13]. More recently, the Local/Equivariant Brouwer Degrees and their infinite dimensional generalizations – the Local/Equivariant Leray-Schauder Degrees – have proven prodigiously effective in obtaining existence results for solutions to a wide class of nonlinear differential equations. Somewhat remarkably, these degree theories are also instrumental in solving classical and equivariant bifurcation problems. The first use of the Leray-Schauder degree for the detection of (local) bifurcation in a parameterized system of nonlinear differential equations is attributed to the seminal work of M.A. Krasnosel’skii, in which sufficient conditions for the existence of a branch of non-trivial solutions are established (cf. [17]). Only a decade later, P. Rabinowitz (cf. [20]) proposed his famous Rabinowitz alternative, which provides sufficient conditions for the unboundedness of a branch of non-trivial solutions. In this paper, we employ equivariant analogues of Krasnosel’skii and Rabinowitz type results to solve an equivariant bifurcation problem. Specifically, our objective is to study the equivariant bifurcation problem for a symmetric system of parameterized nonlinear Laplace equations subject to Dirichlet boundary conditions. The bifurcation and global behaviour of solutions for such systems has been extensively studied, due to their significance in applied mathematics and physics. In the case that the nonlinearity arises as the gradient of some potential function, the bifurcation problem can be studied as a parameterized variational problem and appropriate topological methods such as Morse theory, index theory, and the equivariant gradient degree (cf. [14]) can be administered (cf. [6, 7, 8, 10, 12, 11, 21, 22, 23, 24, 25, 26]). However, these approaches are not applicable for systems which lack variational structure. The degree theoretic methods used in this paper impose no variational requirements on the equations and, as such, are broadly applicable to a wider class of problems. Consider the following parameterized family of symmetric Laplace equations:
@@
The methods used in this paper are inspired by [2], where an existence result was obtained for a similar sublinear elliptic system using equivariant degree theory. For a more thorough exposition of these topics, we direct readers to the recent monograph [4]. The equivariant degree is intimately connected with the classical Brouwer degree. Although its application is relatively simple, technical difficulties arise when dealing with algebraic computations related to unfamiliar group structures. Many of these issues can be resolved with usage of the G.A.P. system, and the G.A.P. package EquiDeg (created by Haopin Wu), which is available online at https://github.com/psistwu/equideg (cf. [28]). In the remainder of this paper, we employ tools from the equivariant degree theory to determine $(i)$ under what conditions branches of non-trivial solutions may bifurcate from a trivial solution, $(ii)$ under what conditions these branches consist only of non-radial solutions and $(iii)$ under what conditions these branches are unbounded. Subsequent sections are organized as follows: In Section 2 the problem (1) is reformulated in an appropriate functional setting. In Section 3 we recall the abstract equivariant bifurcation results, including the equivariant analogues of the classical Krasnosel’skii and Rabinowitz theorems. In Section 4, we apply the equivariant degree theory methods to establish local and global bifurcation results for (1). Finally, in Section 5 we present a motivating example of vibrating membranes where our main results, Theorem 4.1 and 4.3, are applied to demonstrate the existence of unbounded branches of non-radial solutions admitting all possible maximal orbit types. For convenience, the Appendices include an explanation of notations used, a summary of the spectral properties of the Laplace operator on the unit disc, and a brief introduction to the Brouwer equivariant degree theory. Acknowledgment: We are deeply grateful to our advisor, Professor Krawcewicz, whose invaluable guidance and unwavering support made this work possible. We would also like to acknowledge Professor Garcia-Azpetia for insightful discussions regarding the estimation of the behaviour of our system near a bifurcation point.
@@ 3. Abstract Local and Global Equivariant Bifurcation
In this section we present a concise exposition, following [4], of an equivariant Brouwer degree method to study symmetric bifurcation problems. Given a compact Lie group $\mathcal{G}$, an isometric Banach $\mathcal{G}$-representation $\mathcal{H}$ and a completely continuous $\mathcal{G}$-equivariant field $\mathcal{F}:\mathbb{R}\times\mathcal{H}\rightarrow\mathcal{H}$, we use the Leray-Schauder Equivariant $\mathcal{G}$-Degree to describe local and global properties of the solution set to the equation,
@@ 3.1. The Local Bifurcation Invariant and Krasnosel’skii’s Theorem
the local bifurcation invariant at $(\lambda_{0},0)$. The reader is referred to [4] for proof that the invariant (13) does not depend on the choice of $\alpha^{\pm}_{0}\in\mathbb{R}$ or radius $\delta>0$, and also for the proof of the following local bifurcation result, which is a consequence of the equivariant version of a classical result of K. Kuratowski (cf. [18], Thm. 3, p. 170).
@@ 3.2. Global Bifurcation and the Rabinowitz Alternative
Notice that the local bifurcation invariant $\omega_{\mathcal{G}}(\alpha_{0})$ at any critical point $(\alpha_{0},0)\in\Lambda$ is well-defined under assumption ($B_{3}$). Moreover, if $\mathcal{U}\subset\mathbb{R}\times\mathscr{H}$ is an open bounded $\mathcal{G}$-invariant set, then its intersection with the critical set is finite. These observations will be important for the statement of the following global bifurcation result, the proof of which can be found in [4].
@@ 4. Local and Global Bifurcation of Non-radial Solutions in (1)
$(\theta,\gamma,\pm 1)u(z)$ :=±$:=\pm\gamma u(e^{i\theta}\cdot z),\quad z\in D,\;\;u\in\mathscr{H};$
@@
Before approaching the question of bifurcation in the equation (1), we must derive a workable formula for the computation of the degree $G\text{\rm-deg}(\mathscr{A}(\alpha),B(\mathscr{H}))$, where $B(\mathscr{H}):=\{u\in\mathscr{H}:\|u\|_{\mathscr{H}}<1\}$ is the open unit ball in $\mathscr{H}$ and $\alpha\in\mathbb{R}$ is any parameter value for which $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism. Assuming that a complete list of the irreducible $\Gamma$-representations $\{\mathcal{V}_{j}\}_{j=1}^{r}$ is made available we denote by $\{\mathcal{V}_{j}^{-}\}_{j=1}^{r}$ the corresponding list of irreducible $\Gamma\times\mathbb{Z}_{2}$-representations, where the superscript is meant to indicate that each irreducible $\Gamma$-representation has been equipped with the antipodal $\mathbb{Z}_{2}$-action in the standard way (cf. [4]). As a $\Gamma$-representation, $V=\mathbb{R}^{k}$ is also a natural $\Gamma\times\mathbb{Z}_{2}$-representation with the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition
@@
$\mathscr{E}_{nm}^{j}\simeq\mathcal{W}_{m}\otimes\mathcal{V}_{j},\;(n,m)\in{\mathbb{N}}\times{\mathbb{N}}\cup\{0\},\;j\in\{1,2,\ldots,r\}.$
@@
$\square$ For any regular point of equation (8) $\alpha\in\mathbb{R}$, the product property of the Leray-Schauder $G$-equivariant degree (cf. Appendix C) permits us to express $G\text{\rm-deg}(\mathscr{A}(\alpha),B(\mathscr{H}))$ in terms of a Burnside ring product of the Leray-Schauder $G$-equivariant degrees of the various restrictions $\mathscr{A}_{n,m}^{j}(\alpha):\mathscr{E}^{j}_{nm}\rightarrow\mathscr{E}^{j}_{nm}$ of the $G$-equivariant linear isomorphism $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ to the $G$-subrepresentations $\mathscr{E}^{j}_{nm}$ on their respective open unit balls $B(\mathscr{E}^{j}_{nm}):=\{u\in\mathscr{E}^{j}_{nm}\;:\;\|u\|_{\mathscr{H}}<1\}$ as follows
@@ 4.1. Computation of the Local Bifurcation Invariant
In order to formulate the main local equivariant bifurcation result, we must first consider some additional properties of the basic degree. For a more thorough exposition of these topics we refer the reader to [4]. Take $s\in\mathbb{N}$ and define the $s$-folding map as the Lie group homomorphism,
$\psi_{s}(e^{i\theta},\gamma,\pm 1)=(e^{sl\theta},\gamma,\pm 1),\quad\psi_{s}(\kappa e^{i\theta},\gamma,\pm 1)=(\kappa e^{is\theta},\gamma,\pm 1).$
@@
An orbit type which is maximal in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ is also maximal in $\Phi_{0}(G;\mathscr{H}_{0}\setminus\{0\})$. Therefore, any $u\in\mathscr{H}\setminus\{0\}$ with an isotropy $G_{u}\leq G$ such that $(G_{u})$ is maximal in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must be radially symmetric. It follows that in order to detect branches of solutions to (7) corresponding to maps which are both non-trivial and non-radial, we must restrict our focus to orbit types which are maximal in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some positive $m\in\mathbb{N}$. Indeed, to demonstrate the existence of a branch of non-radial solutions bifurcating from some isolated critical point $(\alpha_{0},0)\in\Lambda$, it is sufficient to show that for some $m>0$ there is an orbit type $(H)\in\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ with $\operatorname{coeff}^{H}(\omega_{G}(\alpha_{0}))\not=0$.In such a case, one can additionally conclude that for some $s\geq 1$ there exists a branch non-radial solutions to (7) with symmetries at least $(^{s}H)$ bifurcating from $(\alpha_{0},0)$.
@@ 5. Motivating Example: Vibrating Membranes with Symmetric Coupling
defined for some fixed $c,d\in\mathbb{R}$ satisfying $c>0$, $d<0$, $4d+c\geq 0$ and where $\zeta:\mathbb{R}\rightarrow\mathbb{R}$ is the sigmoid function
@@
Let us specify a particular bifurcation problem associated with system (42) by choosing the following values for the constants $a,c,d\in\mathbb{R}$
@@
2
3LoadPackage ("EquiDeg");
4# create groups O(2)xS4xZ2
5o2:=OrthogonalGroupOverReal(2);
6s4:=SymmetricGroup(4);
7z2:=pCyclicGroup(2);
8# generate S4xZ2
9g1:=DirectProduct(s4,z2);
10# set names for subgroup conjugacy classes in S4xZ2
11SetCCSsAbbrv(g1, [ "Z1", "Z2", "D1z","D1","Z2m", "Z1p",
12"Z3", "V4", "D2z", "Z4", "D2", "D1p","D2d",
13"V4m", "D2p", "Z4d", "D3", "D3z", "Z3p",
14"V4p", "D4z", "D4d", "Z4p", "D4", "D4z",
@@
16# generate O(2)xS4xZ2
17G:=DirectProduct(o2,g1);
18ccs:=ConjugacyClassesSubgroups(G);
19# find Maximal Orbit Types and Basic Degrees in first O(2)-isotypic component
20irrs := Irr(G);
21# The G-istoypic components 0,2,3 are indexed in G.A.P as 3,5,7
22for i in [ 3,5,7 ] do

### v3

@@ 1. Introduction
We draw our motivation from natural phenomena that are modeled by parameterized systems of autonomous Partial Differential Equations (PDEs), where the possible states of each phenomenon are represented by solutions of the associated system of equations at corresponding parameter values. These solutions can be categorized as trivial or non-trivial. Trivial solutions are so-named because they persist for all parameter values and their existence is straightforward to determine. In contrast, non-trivial solutions may only exist for a particular range of parameter values and often exhibit exotic properties that enhance our understanding of the model. The classical bifurcation problem is concerned with the existence and global behaviour of branches of non-trivial solutions emerging from their trivial counterparts. Phenomena which admit the symmetries of a certain group $G$ (represented by the $G$-equivariance of the associated equations) may be studied as equivariant bifurcation problems, with additional consideration placed on the symmetric properties of these branches (cf. [1, 3, 9, 11, 19, 21, 22, 23, 24, 25]). The application of topological methods to the study of differential equations, as with most tools in the arsenal of the nonlinear analyst, traces back to M. Poincare [13]. More recently, the Local/Equivariant Brouwer Degrees and their infinite dimensional generalizations – the Local/Equivariant Leray-Schauder Degrees – have proven prodigiously effective in obtaining existence results for solutions to a wide class of nonlinear differential equations. Somewhat remarkably, these degree theories are also instrumental in solving classical and equivariant bifurcation problems. The first use of the Leray-Schauder degree for the detection of (local) bifurcation in a parameterized system of nonlinear differential equations is attributed to the seminal work of M.A. Krasnosel’skii, in which sufficient conditions for the existence of a branch of non-trivial solutions are established (cf. [17]). Only a decade later, P. Rabinowitz (cf. [20]) proposed his famous Rabinowitz alternative, which provides sufficient conditions for the unboundedness of a branch of non-trivial solutions. In this paper, we employ equivariant analogues of Krasnosel’skii and Rabinowitz type results to solve an equivariant bifurcation problem. Specifically, our objective is to study the equivariant bifurcation problem for a symmetric system of parameterized nonlinear Laplace equations subject to Dirichlet boundary conditions. The bifurcation and global behaviour of solutions for such systems has been extensively studied, due to their significance in applied mathematics and physics. In the case that the nonlinearity arises as the gradient of some potential function, the bifurcation problem can be studied as a parameterized variational problem and appropriate topological methods such as Morse theory, index theory, and the equivariant gradient degree (cf. [14]) can be administered (cf. [6, 7, 8, 10, 12, 11, 21, 22, 23, 24, 25, 26]). However, these approaches are not applicable for systems which lack variational structure. The degree theoretic methods used in this paper impose no variational requirements on the equations and, as such, are broadly applicable to a wider class of problems. Consider the following parameterized family of symmetric Laplace equations:
@@
The methods used in this paper are inspired by [2], where an existence result was obtained for a similar sublinear elliptic system using equivariant degree theory. For a more thorough exposition of these topics, we direct readers to the recent monograph [4] or, alternatively, the older text [5]. The equivariant degree is intimately connected with the classical Brouwer degree. Although its application is relatively simple, technical difficulties arise when dealing with algebraic computations related to unfamiliar group structures. Many of these issues can be resolved with usage of the G.A.P. system, and the G.A.P. package EquiDeg (created by Haopin Wu), which is available online at https://github.com/psistwu/equideg (cf. [28]). In the remainder of this paper, we employ tools from the equivariant degree theory to determine $(i)$ under what conditions branches of non-trivial solutions may bifurcate from a trivial solution, $(ii)$ under what conditions these branches consist only of non-radial solutions and $(iii)$ under what conditions these branches are unbounded. Subsequent sections are organized as follows: In Section 2 the problem (1) is reformulated in an appropriate functional setting. In Section 3 we recall the abstract equivariant bifurcation results, including the equivariant analogues of the classical Krasnosel’skii and Rabinowitz theorems. In Section 4, we apply the equivariant degree theory methods to establish local and global bifurcation results for (1). Finally, in Section 5 we present a motivating example of vibrating membranes where our main results, Theorem 4.1 and 4.3, are applied to demonstrate the existence of unbounded branches of non-radial solutions admitting all possible maximal orbit types. For convenience, the Appendices include an explanation of notations used, a summary of the spectral properties of the Laplace operator on the unit disc, and a brief introduction to the Brouwer equivariant degree theory. Acknowledgment: We are deeply grateful to our advisor, Professor Krawcewicz, whose invaluable guidance and unwavering support made this work possible. We would also like to acknowledge Professor Garcia-Azpetia for insightful discussions regarding the estimation of the behaviour of our system near a bifurcation point.
@@ 3. Abstract Local and Global Equivariant Bifurcation
In this section we present a concise exposition, following [4] and [5], of an equivariant Brouwer degree method to study symmetric bifurcation problems. Given a compact Lie group $\mathcal{G}$, an isometric Banach $\mathcal{G}$-representation $\mathcal{H}$ and a completely continuous $\mathcal{G}$-equivariant field $\mathcal{F}:\mathbb{R}\times\mathcal{H}\rightarrow\mathcal{H}$, we use the Leray-Schauder Equivariant $\mathcal{G}$-Degree to describe local and global properties of the solution set to the equation,
@@ 3.1. The Local Bifurcation Invariant and Krasnosel’skii’s Theorem
the local bifurcation invariant at $(\lambda_{0},0)$. The reader is referred to [4] or [5] for proof that the invariant (13) does not depend on the choice of $\alpha^{\pm}_{0}\in\mathbb{R}$ or radius $\delta>0$, and also for the proof of the following local bifurcation result, which is a consequence of the equivariant version of a classical result of K. Kuratowski (cf. [18], Thm. 3, p. 170).
@@ 3.2. Global Bifurcation and the Rabinowitz Alternative
Notice that the local bifurcation invariant $\omega_{\mathcal{G}}(\alpha_{0})$ at any critical point $(\alpha_{0},0)\in\Lambda$ is well-defined under assumption ($B_{3}$). Moreover, if $\mathcal{U}\subset\mathbb{R}\times\mathscr{H}$ is an open bounded $\mathcal{G}$-invariant set, then its intersection with the critical set is finite. These observations will be important for the statement of the following global bifurcation result, the proof of which can be found in [4], [5].
@@ 4. Local and Global Bifurcation of Non-radial Solutions in (1)
(14) $(\theta,\gamma,\pm 1)u(z)$ :=±$:=\pm\gamma u(e^{i\theta}\cdot z),\quad z\in D,\;\;u\in\mathscr{H};$
@@
Before approaching the question of bifurcation in the equation (1), we must derive a workable formula for the computation of the degree $G\text{\rm-deg}(\mathscr{A}(\alpha),B(\mathscr{H}))$, where $B(\mathscr{H}):=\{u\in\mathscr{H}:\|u\|_{\mathscr{H}}<1\}$ is the open unit ball in $\mathscr{H}$ and $\alpha\in\mathbb{R}$ is any parameter value for which $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism. Assuming that a complete list of the irreducible $\Gamma$-representations $\{\mathcal{V}_{j}\}_{j=1}^{r}$ is made available we denote by $\{\mathcal{V}_{j}^{-}\}_{j=1}^{r}$ the corresponding list of irreducible $\Gamma\times\mathbb{Z}_{2}$-representations, where the superscript is meant to indicate that each irreducible $\Gamma$-representation has been equipped with the antipodal $\mathbb{Z}_{2}$-action in the standard way (cf. [4], [5]). As a $\Gamma$-representation, $V=\mathbb{R}^{k}$ is also a natural $\Gamma\times\mathbb{Z}_{2}$-representation with the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition
@@
$\mathscr{E}_{nm}^{j}\simeq\mathcal{W}_{m}\otimes\mathcal{V}_{j}^{-},\;(n,m)\in{\mathbb{N}}\times{\mathbb{N}}\cup\{0\},\;j\in\{1,2,\ldots,r\}.$
@@
$\square$ The product property of the Leray-Schauder $G$-equivariant degree (cf. Appendix C) permits us to express $G\text{\rm-deg}(\mathscr{A}(\alpha),B(\mathscr{H}))$, at any regular point $\alpha\in\mathbb{R}$ of equation (8), in terms of a Burnside ring product of the Leray-Schauder $G$-equivariant degrees of the various restrictions $\mathscr{A}_{n,m}^{j}(\alpha):\mathscr{E}^{j}_{nm}\rightarrow\mathscr{E}^{j}_{nm}$ of the $G$-equivariant linear isomorphism $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ to the $G$-subrepresentations $\mathscr{E}^{j}_{nm}$ on their respective open unit balls $B(\mathscr{E}^{j}_{nm}):=\{u\in\mathscr{E}^{j}_{nm}\;:\;\|u\|_{\mathscr{H}}<1\}$ as follows
@@ 4.1. Computation of the Local Bifurcation Invariant
In order to formulate the main local equivariant bifurcation result, we must first consider some additional properties of the basic degree. For a more thorough exposition of these topics we refer the reader to [4], [5]. Take $s\in\mathbb{N}$ and define the $s$-folding map as the Lie group homomorphism,
$\psi_{s}(e^{i\theta},\gamma,\pm 1)=(e^{s\theta},\gamma,\pm 1),\quad\psi_{s}(\kappa e^{i\theta},\gamma,\pm 1)=(\kappa e^{is\theta},\gamma,\pm 1).$
@@
An orbit type which is maximal in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ is also maximal in $\Phi_{0}(G;\mathscr{H}_{0}\setminus\{0\})$. Therefore, any $u\in\mathscr{H}\setminus\{0\}$ with an isotropy $G_{u}\leq G$ such that $(G_{u})$ is maximal in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must be radially symmetric. In order to detect branches of solutions to (7) corresponding to maps which are both non-trivial and non-radial, we must restrict our focus to orbit types which are maximal in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some positive $m\in\mathbb{N}$. Indeed, to demonstrate the existence of a branch of non-radial solutions bifurcating from some isolated critical point $(\alpha_{0},0)\in\Lambda$, it is sufficient to show that for some $m>0$ there is an orbit type $(H)\in\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ with $\operatorname{coeff}^{H}(\omega_{G}(\alpha_{0}))\not=0$. In such a case, one can additionally conclude that for some $s\geq 1$ there exists a branch non-radial solutions to (7) with symmetries at least $(^{s}H)$ bifurcating from $(\alpha_{0},0)$.
@@ 5. Motivating Example: Vibrating Membranes with Symmetric Coupling
defined for some fixed $c,d\in\mathbb{R}$ satisfying the conditions
  1. ($E_{4}$)
    $c>0$, $d<0$, $4d+c\geq 0$,
and where $\zeta:\mathbb{R}\rightarrow\mathbb{R}$ is the sigmoid function
@@
We are now in a position to apply the local equivariant bifurcation results derived in this paper, namely Theorem 4.1 and Corollary 4.2, to the example system (43).
Proposition 5.1.
Under assumptions ($E_{1}$)($E_{4}$), every critical point $\alpha_{n,m,j}$ of (43) is a branching point for a branch of non-trivial solutions with symmetries at least $({}^{s}H)$ for some orbit type of maximal kind $(H)\in\mathfrak{M}_{1,j}$ and some folding $s\geq m$.
Proof.
Let $\alpha_{n,m,j}$ be a critical point of (43) with a deleted regular neighborhood $\alpha_{n,m,j}^{-}<\alpha_{n,m,j}<\alpha_{n,m,j}^{+}$ on which $\mathscr{A}(\alpha)$ is an isomorphism and choose any orbit type of maximal kind $(H)\in\mathfrak{M}_{1,j}$ with $2\nmid\dim\mathcal{V}_{1,j}^{H}$. The unique identification of $(n,m,j)\in\Sigma$ with $\alpha_{n,m,j}$ (cf. (50)) together with the strict monotonicity of the eigenvalues (49) imply the set difference $\Sigma(\alpha_{n,m,j}^{+})\setminus\Sigma(\alpha_{n,m,j}^{-})=\{(n,m,j)\}$. Therefore, the numbers $\mathfrak{n}^{s}(\alpha_{n,m,j}^{\pm},H)$ agree for all $s\neq m$ and are consecutive in the case of $s=m$ and the result follows from Corollary (4.2). $\square$ In order to demonstrate how Theorem 4.3, the main global equivariant bifurcation result of this paper, can be applied to a system of the form (43), we introduce the following simplifying assumption:
  1. ($E_{5}$)
    there exists an isotypic index $j\in\{0,2,3\}$, an odd number $m_{0}\in 2{\mathbb{N}}-1$ and a pair of numbers $n_{l},n_{u}\in{\mathbb{N}}$ with $n_{l}\leq n_{u}$ and $2\mid n_{u}-n_{l}$ such that the Bessel root $s_{nm}$ lies:
    1. (i)
      within the codomain of $\mu_{j}$ if $m=m_{0}$ and $n_{l}\leq n\leq n_{u}$;
    2. (ii)
      outside the codomain of $\mu_{j}$ if either, $m=m_{0}$ and $n<n_{l}$ or $n>n_{u}$, or if $m>m_{0}$ for any $n\in{\mathbb{N}}$.
Condition ($E_{5}$) guarantees that an odd number of critical points $(\alpha,0)\in\Lambda^{\bm{K}}$ will obtain $\mathfrak{s}(\alpha,H)=\bar{\mathfrak{s}}(H)$ for every orbit type of maximal kind $(H)\in\mathfrak{M}_{1,j}$ with $2\nmid\dim\mathcal{V}_{1,j}^{H}$.
Proposition 5.2.
Under assumptions ($E_{1}$)($E_{5}$), the trivial solution of (43) undergoes a global bifurcation of non-radial solutions with symmetries at least $({}^{s}H)$ at every critical point $\alpha_{n,m_{0},j}$ with $n_{l}\leq n\leq n_{u}$ for every orbit type of maximal kind $(H)\in\mathfrak{M}_{1,j}$ with $2\nmid\dim\mathcal{V}_{1,j}^{H}$.
Proof.
Choose $(H)\in\mathfrak{M}_{1,j}$ with $2\nmid\dim\mathcal{V}_{1,j}^{H}$ and notice that, since $\bar{\mathfrak{s}}(H)=m_{0}$ and $\alpha_{n,m_{0},j}\in\mathfrak{J}(H)$ for every $n_{l}\leq n\leq n_{u}$, one has $2\nmid|\mathfrak{J}(H)|$ and the result follows from Theorem 4.3. $\square$ So that Propositions 5.1 and 5.2 can be verified empirically, let us specify a particular bifurcation problem associated with system (43) by choosing the following values for the constants $a,c,d\in\mathbb{R}$
@@
2LoadPackage ("EquiDeg");
3# create groups O(2)xS4xZ2
4o2:=OrthogonalGroupOverReal(2);
5s4:=SymmetricGroup(4);
6z2:=pCyclicGroup(2);
7# generate S4xZ2
8g1:=DirectProduct(s4,z2);
9# set names for subgroup conjugacy classes in S4xZ2
10SetCCSsAbbrv(g1, [ "Z1", "Z2", "D1z","D1","Z2m", "Z1p",
11"Z3", "V4", "D2z", "Z4", "D2", "D1p","D2d",
12"V4m", "D2p", "Z4d", "D3", "D3z", "Z3p",
13"V4p", "D4z", "D4d", "Z4p", "D4", "D4z",
@@
15# generate O(2)xS4xZ2
16G:=DirectProduct(o2,g1);
17ccs:=ConjugacyClassesSubgroups(G);
18# find Maximal Orbit Types and Basic Degrees in first O(2)-isotypic component
19irrs := Irr(G);
20# The G-istoypic components 0,2,3 are indexed in G.A.P as 3,5,7
21for i in [ 3,5,7 ] do